OpenStudy (anonymous):

Divide. Write your answer in lowest terms 3/22 / (- 3/8)

OpenStudy (anonymous):

@cupcakeworld

OpenStudy (timaashorty):

\[\frac{ 3 }{ 22} \div (- \frac{ 3 }{ 8 })= -\frac{ 4 }{ 11 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 22} \div \frac{ -3 }{ 8 }\]

OpenStudy (anonymous):

\[\frac{ 3}{ 22 } *\frac{ 8 }{ -3}\]

OpenStudy (anonymous):

3 will get cancelled =-8/22=-4/11